Circular No. 5788

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993N IN UGC 5695 A. V. Filippenko and T. Matheson, University of California at Berke-

ley, report that a relatively noisy spectrum (range 420-700 nm, resolution 0.6 nm) of SN 1993N (cf. IAUC 5784) obtained on May 11 UT with the Shane 3-m reflector at Lick Observatory reveals that the object is a peculiar type-II supernova similar to SN 1988Z. The continuum is featureless, and Hα consists of three emission components: narrow (essentially unresolved), intermediate (FWZI ∼ 2700 km/s), and broad (FWZI ∼ 19 000 km/s). No Hα absorption is visible. SN 1993N is superposed on an H II region in the parent galaxy; the derived redshift is 2800 km/s.

HD 98800 S. Rucinski, Space Astrophysics Laboratory, Institute for Space and

Terrestrial Science, North York, ON, writes: “The remarkable K5 V star HD 98800, which shows the infrared excess attributable to zodiacal dust with mass six orders of magnitude greater than in the solar system (Zuckerman and Becklin 1992, Ap.J. 406, L25) has been detected as a strong 0.8-mm source with the James Clerk Maxwell Telescope (+ UKT14) at Mauna Kea on Apr. 25 under good observing conditions. The flux density at 0.8 mm is 102 ± 10 mJy.”

SUPERNOVA 1993J IN NGC 3031 Further visual magnitude estimates: Mar. 26.995 UT, [17: (O. Schnei-

der, Bielefeld, Germany; Kodak 2415 film); Apr. 28.90, 11.9 (A. Mizser, Budapest, Hungary); May 1.01, 11.9 (B. H. Granslo, Fjellhamar, Norway); 4.92, 12.2 (P. Schmeer, Bischmisheim, Germany); 5.94, 12.3 (H. Dahle, Oslo, Norway); 7.88, 12.4 (Schmeer); 11.11, 12.6 (W. G. Dillon, Missouri City, TX); 11.93, 12.3 (J. D. Shanklin, Cambridge, England).

PERIODIC COMET VÄISÄLÄ 1 (1992u) Total visual magnitude estimates (cf. IAUC 5699): Mar. 13.31 UT,

12.5 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); Apr. 14.54, 13.8 (A. Nakamura, Kuma, Japan, 0.60-m reflector); 17.28, 13.2 (Morris); 26.95, 14.2 (H. Mikuž, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + V filter + ST-6 CCD).

1993 May 13 Daniel W. E. Green